OpenStudy (anonymous):

Integrate e^x/(1+e^(2x)) using substitution?

OpenStudy (anonymous):

looks like \[\int\frac{u}{1+u^2}du\]

OpenStudy (anonymous):

Ahh, that was pretty easy! Been at this a little too long I think. Thanks!

OpenStudy (anonymous):

anti derivative is \[\frac{1}{2}\ln(u^2+1)\]

OpenStudy (anonymous):

i mean you still have to make the substitution yes?

OpenStudy (anonymous):

well, yeah, but I was having a hard time figuring out the form. From there it would be inverse tan(e^x)+C, correct?

OpenStudy (anonymous):

wait wait

OpenStudy (anonymous):

i steered you wrong. you have it. right exactly

OpenStudy (anonymous):

\[u=e^x\] etc. you get \[\tan^{-1}(u)=\tan^{-1}(e^x)\]

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (anonymous):

latex all messed up

OpenStudy (anonymous):

yw